OpenStudy (vane11):

A set of four cards consists of two red cards and two black cards. The cards are shuffled thoroughly, and I am dealt two cards. I found the number of red cards X in these two cards. The random variable X has which of the following probability distributions?

jimthompson5910 (jim_thompson5910):

were you able to get anywhere?

OpenStudy (vane11):

with house question? I think I got... I can't remember what I got, but I did get it right

jimthompson5910 (jim_thompson5910):

you don't have your work with you?

OpenStudy (vane11):

.1075 :)

jimthompson5910 (jim_thompson5910):

how did you get that

OpenStudy (vane11):

Honestly, I got the answer elsewhere, I have a formula for it, but I'm not sure it is correct since I haven't studied yet, I'm just getting all the work in today since it's due, then going over it tomorrow for the test

jimthompson5910 (jim_thompson5910):

ok let's try to work through it and we'll see how things go, sound good?

OpenStudy (vane11):

Sure, the previous one or this one?

jimthompson5910 (jim_thompson5910):

let's do this one first, I don't remember the previous one too well (but we can go back to that one later)

jimthompson5910 (jim_thompson5910):

anyways, what they are looking for here is a table that shows all the possible outcomes of any event and the probabilities of each event possible

jimthompson5910 (jim_thompson5910):

We have the 3 possibilities when it comes to drawing red cards a) Draw 0 red cards ... ie X = 0 b) Draw 1 red card ... X = 1 c) Draw 2 red cards ... X = 2

jimthompson5910 (jim_thompson5910):

what is the probability of drawing 0 red cards?

jimthompson5910 (jim_thompson5910):

put another way, what is the probability of drawing 2 black cards?

OpenStudy (vane11):

50%?

jimthompson5910 (jim_thompson5910):

not quite

OpenStudy (vane11):

would it be 1/12?

OpenStudy (vane11):

no no

OpenStudy (vane11):

1/6

jimthompson5910 (jim_thompson5910):

we have 4 cards and we're picking 2 of them so we have 4 C 2 = 4*3/2 = 6 possible ways to pick 2 cards only one of these ways has 2 black cards

jimthompson5910 (jim_thompson5910):

so yes, it's 1/6

jimthompson5910 (jim_thompson5910):

so using probability notation, we would say P(X = 0) = 1/6 this means that the probability of picking 0 red cards is 1/6

jimthompson5910 (jim_thompson5910):

what is the probability of picking 1 red card?

OpenStudy (vane11):

i hate looking stupid but 1/3?

jimthompson5910 (jim_thompson5910):

you have 12 ways of picking 2 cards and the 12 ways are R1R2, R1B1, R1B2, R1R2, R2B1, R2B2, R1B1, R2B1, B1B2, R1B2, R2B2, B1B2

jimthompson5910 (jim_thompson5910):

oh sry, I double counted, let me redo

jimthompson5910 (jim_thompson5910):

there are 6 ways to pick 2 cards and the 6 ways are R1R2, R1B1, R1B2, R2B1, R2B2, B1B2

jimthompson5910 (jim_thompson5910):

of these 6 ways, there are 4 ways (shown below) where you can pick exactly 1 red card R1B1, R1B2, R2B1, R2B2

jimthompson5910 (jim_thompson5910):

so the probability of picking exactly one red card is 4/6 = 2/3

jimthompson5910 (jim_thompson5910):

the probability of picking exactly 2 red cards (look at the same list) is 1/6

OpenStudy (vane11):

ok...

jimthompson5910 (jim_thompson5910):

so the probability distribution will look like this